Evaluate the following integral by first converting to an integral in polar coordinates.

0

2


4

y
2


4

y
2
x
2
d
x
d
y
∫−20∫−4−y24−y2x2dxdy
1
Expert's answer
2020-04-22T19:35:39-0400
"\displaystyle\int_{-2}^0\displaystyle\int_{-\sqrt{4-y^2}}^{\sqrt{4-y^2}}x^2dxdy"

"x=r\cos\theta, y=r\sin\theta"

"0\leq r\leq 2, -\pi\leq \theta\leq 0"


"\displaystyle\int_{-2}^0\displaystyle\int_{-\sqrt{4-y^2}}^{\sqrt{4-y^2}}x^2dxdy="

"=\displaystyle\int_{-\pi}^0\displaystyle\int_{0}^{2}rr^2\cos^2\theta drd\theta="

"=\displaystyle\int_{-\pi}^0\big[{r^4 \over 4}\big]\begin{array}{cc}
 2 \\
 0
\end{array}\cos^2\theta d\theta="

"=2\displaystyle\int_{-\pi}^0
(1+\cos(2\theta)) d\theta="

"=\big[2\theta+\sin(2\theta)\big]\begin{array}{cc}
 0\\
 -\pi
\end{array}=2\pi"